OpenStudy (anonymous):

a circular field has a perimeter of 660 m.A plot in the shape of a square,having its vertices on the circumference,is marked in the field.Calculate the area of the square plot

OpenStudy (anonymous):

hint (diagonal of the sq. plot is the Dia)

OpenStudy (anonymous):

i'm confused:(

OpenStudy (anonymous):

do u know the formula for perimeter

OpenStudy (anonymous):

2*pi*r

OpenStudy (anonymous):

u can figure out R

OpenStudy (anonymous):

2100cm?

OpenStudy (anonymous):

210cm*

OpenStudy (anonymous):

2R= dia of the circle

OpenStudy (anonymous):

p=2*pi*r 660=2*22/7*r 660*7/22=r 210 m=r

OpenStudy (anonymous):

try drawing the figure first then u will get a better idea of what is happening

OpenStudy (anonymous):

dia ofcircle = 2^1/2(a) a = side of sq.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

210 SQ CM IS A?

OpenStudy (anonymous):

210 is radius dia will be twice of that=420cm a=(2*210)/2^1/2==210*2^1/2 ar of the plot ==2*210*210

OpenStudy (anonymous):

8820?

OpenStudy (anonymous):

88200sq cm

OpenStudy (anonymous):

more help?more questions?